OpenStudy (anonymous):

Suppose the minute hand on a clock is 15 inches long and the hour hand is 10 inches long. How fast is the distance between the tips of the hands changing at 3:35 PM?

OpenStudy (anonymous):

It's a rates of change problem, and I can't seem to set it up.

OpenStudy (anonymous):

Hmmm, first model the equations in terms of \(\theta\).

OpenStudy (goformit100):

Hmmm, first model the equations in terms of θ.

OpenStudy (anonymous):

Consider that \(x=r\sin\theta,y=r\sin\theta\)

OpenStudy (goformit100):

@Opcode

OpenStudy (anonymous):

@theonechewbaca Make sense?

OpenStudy (anonymous):

is that the formula for the sector that x is in?

OpenStudy (anonymous):

\(x=r\cos\theta,y=r\sin\theta\)

OpenStudy (anonymous):

yeah that seems to be formula for the sector of the circle, so where would I go from there? I have the lengths of each of the hands (x and y), and also the position they are at.

OpenStudy (anonymous):

\[ \frac{dx}{dt} =\frac{d}{dt} r\cos\theta = -r\sin\theta \frac{d\theta}{dt} \]

OpenStudy (anonymous):

This is just the chain rule.

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

We know \[ z^2 = x^2+y^2 \]Differentiating with respect to \(t\) gets us: \[ 2z\frac{dz}{dt}=2x\frac{dx}{dt}+2y\frac{dy}{dt} \]

OpenStudy (anonymous):

In this case \(z\) is our distance, and they're asking for \(dz/dt\)

OpenStudy (anonymous):

Makes sense, and I know what x and y are equal to

OpenStudy (anonymous):

So, I've told you how to find \(dx/dt, dy/dt, x, y,z\) You should be able to do the rest yourself.

OpenStudy (anonymous):

Again \(x=r\cos\theta, y=r\sin\theta\) Where \(r\) is the radius (length of hand) and \(\theta\) is the angle

OpenStudy (anonymous):

and I find the theta by plugging it into radian form?

OpenStudy (anonymous):

So in this case, if its 3:35, the hour hand would be at like -pi/6, and the minute hand would be at -2pi/3?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

Now this is 20 minutes away from 15 minutes.

OpenStudy (anonymous):

right, but at 35 minutes the hour hand would be between 3 and 4 right?

OpenStudy (anonymous):

So the angle is \(-20/60 \times 2\pi\)

OpenStudy (anonymous):

Hmmm, that's true...

OpenStudy (anonymous):

Umm, after 35 minutes, you have 35/60 hours, and (35/60)/12 rotations.

OpenStudy (anonymous):

and at 35 minutes, theta would just be 4pi/3 cause its at the 240 degrees

OpenStudy (anonymous):

*5pi/3

OpenStudy (anonymous):

Ummm, \(-2\pi/3 = 4\pi/3\)

OpenStudy (anonymous):

so now that we've figured out the theta angles for each situation, how do i plug that into the equation to figure out z? I'm thinking that i should put in x=rcos(theta) into the derivative of pythagoreans that we figured out earlier. Do the same for y, and then plug in dx/dt and dy/dt into the same equation and solve for z? But then how do I figure out dz/dt?

OpenStudy (anonymous):

You find z once you have found x and y.

OpenStudy (anonymous):

yeah plug that into the x^2 + y^2 = z^2 equation?

OpenStudy (anonymous):

Then \(dz/dt\) will remain... that is the thing you're trying to find.

OpenStudy (anonymous):

awesome! thanks for all the help, im gonna crash right now but im gonna work on it when im up in the morning. I'll tag you in my final post, and if you dont mind id like you to let me know if i got it right or not

OpenStudy (anonymous):

how am i supposed to figure out d(theta)/dt? This is what I have so far, \[2rcos(\theta)\times(-rsin(\theta))\times [d(\theta)\div(dt)]+2rsin(\theta)\times rcos(\theta) \times (d(\theta)/dt)\]All over 2r. Now all I need to figure out is what d(theta)/dt could be

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Remember \(d\theta/dt\) is just how face the hand moves round the clock!

OpenStudy (anonymous):

For example, for the minute hand it does 1 revelation for every 60 seconds, so it is going to be \(-2\pi/60\)

OpenStudy (anonymous):

The angles are negative because clockwise is backwards.

OpenStudy (anonymous):

@wio, But wait, there are two hands in this equation and only one r, shouldn't there be a R and a r, (hour hand and minute hand, respectively) in the equation?

OpenStudy (anonymous):

Yes, there are two Rs

OpenStudy (anonymous):

Alright let's look at this problem from another angle, using the law of cosines, we could say that the length of the distance between the two hands is equal to \[c^2 = m^2 + h^2 - 2mhcos(\theta)\]

OpenStudy (anonymous):

If we plug in the values for the minute hands and the hour hands, we could get a number, then take the derivative of the whole equation right?

OpenStudy (anonymous):

We don't need to change anything just because the Rs are different.

OpenStudy (anonymous):

I'm not understanding, could you write out the equation?

OpenStudy (anonymous):

I'm really busy right now. Just do whatever you think will work.